S. Lee Enterprises produces two models of lawn tractor: SL1 and SL2. The models are both produced in the company’s Louisville factory and go through identical assembly operations. The difference is in the quality (hence, cost) of the parts used. The following data are available for November, the latest month for which information is available:

SL1 SL2 Total
Number assembled 1,650 2,150 3,800
Materials cost per tractor $ 1,250 $ 1,750
Other costs:
Direct labor $ 1,550,000
Supplies 550,000
Other overhead 2,840,000


Required:

S. Lee Enterprises uses operations costing and assigns conversion costs based on the number of units assembled. Compute the cost of each model assembled in November.

S. Lee Enterprises produces two models of lawn tractor: SL1 and SL2. The models are both produced in the company’s Louisville factory and go through identical assembly operations. The difference is in the quality (hence, cost) of the parts used. The following data are available for November, the latest month for which information is available:

SL1 SL2 Total
Number assembled 1,650 2,150 3,800
Materials cost per tractor $ 1,250 $ 1,750
Other costs:
Direct labor $ 1,550,000
Supplies 550,000
Other overhead 2,840,000


Required:

S. Lee Enterprises uses operations costing and assigns conversion costs based on the number of units assembled. Compute the cost of each model assembled in November.

operation cost    sl1 ?    sl2?    total    ?
materials cost    sl1    ?    sl2? total ?
total cost sl1 ?    sl2? total ?
number of units sl1    ? sl2 ? total    ?
unit cost    sl1 ? sl2 ?    total    ?

answers are required for all sl1    sl2    and total for each one
Answer #1
Computation of total material cost
Particulars SL1 SL2 Total
No of assembled (1) 1650 2150 3800
material cost per tractor (2) 1250 1750
total material cost (1)*(2) 2062500 3762500 5825000
Conversion cost per tractor = total conversion costs/no of unit
total conversion cost = direct labor+supplies+other overhead
total conversion cost = 1550000+550000+2840000
total conversion cost = 4940000
Conversion cost per tractor = 4940000/3800 =1300
total operational cost
SL1 = 1650*1300 = 2145000
SL2 = 2150*1300 = 2795000
Computation of cost of each model assembled in november
Particulars SL1 SL2
Operational cost 2145000 2795000 4940000
material cost 2062500 3762500 5825000
total cost (A) 4207500 6557500 10765000
no of units (B) 1650 2150 3800
unit cost (A)/(B) 2550 3050 5600